# location=local]refs.bib

### v1

@@ II Related Work
Type
Dataset
Data
type
Year # Labels
# Annotations
per frame
# Object
classes
Dynamic
objectsa
Tasksb
@@
2D B3DO [1] Real 2011 849 frames 2$\sim$5 50+ × × × ×
NYU-Depth v2 [2] Real 2012 1,449 frames 30$\sim$40 894 × × × ×
SUN3D [3] Real 2013 8 scans 10$\sim$15 - × × ×
Stanford 2D-3D-S [4] Real 2017 70,496 frames 10$\sim$15 13 × × × ×
SceneNet RGB-D [5] Synthetic 2017 5M frames 20$\sim$30 255 × × ×
3D SUN RGB-D [6] Real 2015 10k frames 20$\sim$30 800 × ×
ScanNet [7] Real 2017 2.5M frames 10$\sim$15 21 × ×
SUN-CG [8] Synthetic 2017 500k frames 5$\sim$15 84 × × ×
Matterport 3D [9] Real 2017 194,400 frames 5$\sim$15 40 × × ×
InteriorNet [10] Synthetic 2018 20M frames 20$\sim$30 158 × × ×
ARKitScenes [11] Real 2022 5,047 scans 5$\sim$10 17 × × × ×
ScanNet++ [12] Real 2023 1,858 scans 20$\sim$30 1,000+ × × ×
@@
Since effective and intensive 3D annotations is hard to acquire, some works label RGB-D images with 2D annotations, which indirectly serve as ground truth for 3D tasks. Berkeley 3D Object Dataset [1] has 2D bounding box annotations on RGB-D images, NYU Depth v2 [2] includes 2D semantic segmentation from short RGB-D videos with 1449 selected frames tagged, SUN3D [3] dataset is composed of 415 RGB-D video sequences in 254 scenes. Stanford 2D-3D-Semantics dataset [4] utilizes the iGibson simulation environment to provide large scale virtual scenes with 2D texture, geometric as well as semantic information. SceneNet RGB-D dataset [5] contains 5000k images and different kinds of 2D annotations.
3D annotation is challenging, yet some works have still made significant contributions. SUN RGB-D [6] contains 10,335 RGB-D images with dense 2D/3D annotations, including 2D polygons, 3D bounding boxes with accurate object orientation and 3D room layout. ScanNet [7] contains a total of 1,513 video sequences with annotations of 3D camera poses, surface reconstruction and semantic segmentation, and partially aligned CAD models. SUN-CG [8] provides 45k virtual scene layouts and 500k ren- dered images with single-view RGB, depth maps and semantic segmentation maps. Matterport 3D [9] contains 194,400 RGB- D images to generate panoramas with surface reconstruction, camera position and 2D/3D semantic segmentation annota- tions. InteriorNet [10] is rendered entirely in virtual home scenes and contains 15k sequences. ARKitScenes [11] improves the resolution of ground truth geometry from laser scans. ScanNet++ [12] is a new dataset that contains 460 high-resolution 3D reconstructions of indoor scenes with dense semantic and instance annotations.
Drawing references from both 2D and 3D annotation works, current datasets lack dynamic objects and pedestrians, which are highly common in real-world scenes [13, 14, 15, 16, 17, 18, 19, 20] and pose sufficient challenges for robot-related research. These datasets commonly exhibit issues such as poor data quality, small annotation volumes, limited label types, noisy annotations, and unreasonable virtual scene layouts. Current research often requires testing, training, and deployment on real hardware in virtual scenarios, which the current datasets cannot fully satisfy.
@@ III-A1 Real Data Acquisition
The image data includes depth and RGB images. We have developed ROS scripts for synchronous acquisition of RGB and depth images. The depth images in the released version have been denoised using the Self-Supervised Deep Depth Denoising method [21]. Due to the inherent instability of Kinect V2, the actual frame rate for capturing RGB and depth images fluctuates between 15-30fps.
@@ III-B2 Real Data Annotation
To annotate the real-world data with 2D and 3D object information, we first preprocessed RGB images and depth maps by applying automated algorithms to predict the 2D&3D positions and bounding boxes of objects, such as Faster-RCNN [22] and Votenet [23]. However, the algorithms for 3D objects detection are not very effective. Subsequently, we performed manual review and verification to correct and refine the annotated bounding boxes, ensuring the accuracy and completeness of the detected objects. For the 2D semantic and instance segmentation task, we similarly employed a semi-automated annotation method, manual review and correction were also conducted by annotators to ensure accurate pixel-wise classification and semantic labeling.
@@ IV Evaluation on THUD Robotic Dataset
Supermarket F-PointNet [24] 7.89 8.92
ImVoteNet [23] 17.49 17.29
DeMF [25] 34.51 38.24
Canteen F-PointNet [24] 18.16 36.67
ImVoteNet [23] 26.72 43.37
DeMF [25] 28.56 45.43
@@
3D object detection methods estimate a 3D bounding box and a pose of each object presented in a scene [26, 27, 28]. To compare the performance difference, three representative indoor 3D object detection algorithms which have good performance on both dynamic and static objects, i.e. F-PointNet [24], ImVoteNet [23] and DeMF [25], are selected. These algorithms were evaluated on the real-world canteen scene and the synthetic supermarket scene to assess their mean Average Precision (mAP) for static and dynamic objects. The experimental results are depicted in Table II. In the supermarket scene, the selected 5 static objects include Chair, Table, Shelf, Cabinet, Fridge. While the 3 dynamic objects encompass Shopping cart, Robot, People. In the canteen scene, the 7 static objects are Chair, Table, Shelf, Bench, Door, Stairs, Stool, and the dynamic object is only People. Moreover, we used the average number of dynamic objects per frame as a metric to assess the level of scene dynamic complexity, the supermarket and canteen scene dynamic complexity is 0.94 and 3.34, respectively.
@@
Supermarket ACNet [29] 3×R50 74.83
RedNet [30] 2×R34 76.92
ESANet [31] 2×R34 78.42
SA-Gate [32] 2×R101 83.19
Canteen ACNett [29] 3×R50 51.85
RedNet [30] 2×R34 59.83
ESANet [31] 2×R34 65.97
SA-Gate [32] 2×R101 58.34
@@ IV-B Semantic Segmentation
Semantic segmentation methods assign each pixel in an image to its corresponding semantic category, focusing on the geometric structure and differentiation of different objects [33, 34, 35]. However, robot indoor semantic segmentation faces several challenges, e.g. complex indoor scenes, occlusions, lighting variations, as well as the diversity of object shapes and sizes. Four RGB-D semantic segmentation algorithms, i.e. ACNet [29], RedNet [30], ESANet [31], and SA-Gate [32], are selected and tested on the aforementioned synthetic supermarket scene and real-world canteen scene. 31 and 19 static and dynamic objects with semantic labels are added to the supermarket and canteen scene, respectively. Objects without labels are set to the void value with corresponding pixels (0, 0, 0).
The accuracy of different semantic segmentation methods is measured by calculating the MIoU (Mean Intersection over Union) during testing and the results are shown in Table III. We can obtain three conclusions. First, the accuracy on real-world canteen scene is much lower than synthetic supermarket scene. Second, the MIoU results are comparable to their original published results, though tested on different datasets. Third, to obtain a quantitative comparison on dynamic objects, we also train and test the ESANet [31] algorithm in the supermarket scene with and without dynamic objects, the MIoU of the test is 78.42% and 79.63% respectively. There is no significant difference, which is coincident with our intuition. In addition, as shown in Fig. 7, different methods still have varying segmentation ability on dynamic object, such as the people and shopping carts.
@@ IV-C Robot Relocalization
Robot relocalization refers to the process of accurately determines and adjusts its position based on the environmental perception information and its own localization capabilities, during its movement [36, 37, 38]. Existing methods can be coarsely classified into global- and local-feature based methods.
We conducted extensive experiments on THUD using a representative global- and local-feature based methods, i.e. NetVlad [39] and FeatLoc [40], respectively. We randomly selected two sequences as testing and the results on both synthetic and real-world scene are visualized in Fig. 6.
In addition, to analyze the effect of the number of dynamic objects (primarily pedestrians) in various scene images, we define a metric to represent a scene’s dynamic complexity, computed as the average number of dynamic pedestrians per frame across the entire dataset. As shown in Fig. 8, with the scene’s dynamic complexity increasing, the methods’ accuracy decreases and FeatLoc [40] has a higher drop in accuracy compared to NetVlad [39], attributed to dynamic objects introducing disturbances and noise in the extraction of scene image features.

### v2

@@ II Related Work
Type Dataset Data type Year # Labels # Annotations per frame # Object classes Dynamic objectsa Tasksb
@@
2D B3DO [B3DO] Real 2011 849 frames 2$\sim$5 50+ × × × ×
NYU-Depth v2 [NYU] Real 2012 1,449 frames 30$\sim$40 894 × × × ×
SUN3D [SUN3D] Real 2013 8 scans 10$\sim$15 - × × ×
Stanford 2D-3D-S [Stanford] Real 2017 70,496 frames 10$\sim$15 13 × × × ×
SceneNet RGB-D [SceneNet] Synthetic 2017 5M frames 20$\sim$30 255 × × ×
3D SUN RGB-D [SUNRGBD] Real 2015 10k frames 20$\sim$30 800 × ×
ScanNet [ScanNet] Real 2017 2.5M frames 10$\sim$15 21 × ×
SUN-CG [SUNCG] Synthetic 2017 500k frames 5$\sim$15 84 × × ×
Matterport 3D [Matterport3D] Real 2017 194,400 frames 5$\sim$15 40 × × ×
InteriorNet [InteriorNet] Synthetic 2018 20M frames 20$\sim$30 158 × × ×
ARKitScenes [ARKitScenes] Real 2022 5,047 scans 5$\sim$10 17 × × × ×
ScanNet++ [ScanNet++] Real 2023 1,858 scans 20$\sim$30 1,000+ × × ×
@@
Since effective and intensive 3D annotations is hard to acquire, some works label RGB-D images with 2D annotations, which indirectly serve as ground truth for 3D tasks. Berkeley 3D Object Dataset [B3DO] has 2D bounding box annotations on RGB-D images, NYU Depth v2 [NYU] includes 2D semantic segmentation from short RGB-D videos with 1449 selected frames tagged, SUN3D [SUN3D] dataset is composed of 415 RGB-D video sequences in 254 scenes. Stanford 2D-3D-Semantics dataset [Stanford] utilizes the iGibson simulation environment to provide large scale virtual scenes with 2D texture, geometric as well as semantic information. SceneNet RGB-D dataset [SceneNet] contains 5000k images and different kinds of 2D annotations.
3D annotation is challenging, yet some works have still made significant contributions. SUN RGB-D [SUNRGBD] contains 10,335 RGB-D images with dense 2D/3D annotations, including 2D polygons, 3D bounding boxes with accurate object orientation and 3D room layout. ScanNet [ScanNet] contains a total of 1,513 video sequences with annotations of 3D camera poses, surface reconstruction and semantic segmentation, and partially aligned CAD models. SUN-CG [SUNCG] provides 45k virtual scene layouts and 500k ren- dered images with single-view RGB, depth maps and semantic segmentation maps. Matterport 3D [Matterport3D] contains 194,400 RGB- D images to generate panoramas with surface reconstruction, camera position and 2D/3D semantic segmentation annota- tions. InteriorNet [InteriorNet] is rendered entirely in virtual home scenes and contains 15k sequences. ARKitScenes [ARKitScenes] improves the resolution of ground truth geometry from laser scans. ScanNet++ [ScanNet++] is a new dataset that contains 460 high-resolution 3D reconstructions of indoor scenes with dense semantic and instance annotations.
Drawing references from both 2D and 3D annotation works, current datasets lack dynamic objects and pedestrians, which are highly common in real-world scenes [dai_scannet_2017, armeni_joint_2017, hua_scenenn_2016, ICL-NUIM, song_suncgsemantic_2016, song_sun_2015, TUM, 12Scenes] and pose sufficient challenges for robot-related research. These datasets commonly exhibit issues such as poor data quality, small annotation volumes, limited label types, noisy annotations, and unreasonable virtual scene layouts. Current research often requires testing, training, and deployment on real hardware in virtual scenarios, which the current datasets cannot fully satisfy.
@@ III-A1 Real Data Acquisition
The image data includes depth and RGB images. We have developed ROS scripts for synchronous acquisition of RGB and depth images. The depth images in the released version have been denoised using the Self-Supervised Deep Depth Denoising method [Sterzentsenko_2019_ICCV]. Due to the inherent instability of Kinect V2, the actual frame rate for capturing RGB and depth images fluctuates between 15-30fps.
@@ III-B2 Real Data Annotation
To annotate the real-world data with 2D and 3D object information, we first preprocessed RGB images and depth maps by applying automated algorithms to predict the 2D&3D positions and bounding boxes of objects, such as Faster-RCNN [FasterRCNN] and Votenet [ImVoteNet]. However, the algorithms for 3D objects detection are not very effective. Subsequently, we performed manual review and verification to correct and refine the annotated bounding boxes, ensuring the accuracy and completeness of the detected objects. For the 2D semantic and instance segmentation task, we similarly employed a semi-automated annotation method, manual review and correction were also conducted by annotators to ensure accurate pixel-wise classification and semantic labeling.
@@ IV Evaluation on THUD Robotic Dataset
Supermarket F-PointNet [F-PointNet] 7.89 8.92
ImVoteNet [ImVoteNet] 17.49 17.29
DeMF [DeMF] 34.51 38.24
Canteen F-PointNet [F-PointNet] 18.16 36.67
ImVoteNet [ImVoteNet] 26.72 43.37
DeMF [DeMF] 28.56 45.43
@@
3D object detection methods estimate a 3D bounding box and a pose of each object presented in a scene [DeformableDETR, wang_tokenfusionmultimodal_2022, wang_frustum_2019]. To compare the performance difference, three representative indoor 3D object detection algorithms which have good performance on both dynamic and static objects, i.e. F-PointNet [F-PointNet], ImVoteNet [ImVoteNet] and DeMF [DeMF], are selected. These algorithms were evaluated on the real-world canteen scene and the synthetic supermarket scene to assess their mean Average Precision (mAP) for static and dynamic objects. The experimental results are depicted in Table II. In the supermarket scene, the selected 5 static objects include Chair, Table, Shelf, Cabinet, Fridge. While the 3 dynamic objects encompass Shopping cart, Robot, People. In the canteen scene, the 7 static objects are Chair, Table, Shelf, Bench, Door, Stairs, Stool, and the dynamic object is only People. Moreover, we used the average number of dynamic objects per frame as a metric to assess the level of scene dynamic complexity, the supermarket and canteen scene dynamic complexity is 0.94 and 3.34, respectively.
@@
Supermarket ACNet [ACNet] 3×R50 74.83
RedNet [RedNet] 2×R34 76.92
ESANet [ESANet] 2×R34 78.42
SA-Gate [SA-Gate] 2×R101 83.19
Canteen ACNett [ACNet] 3×R50 51.85
RedNet [RedNet] 2×R34 59.83
ESANet [ESANet] 2×R34 65.97
SA-Gate [SA-Gate] 2×R101 58.34
@@ IV-B Semantic Segmentation
Semantic segmentation methods assign each pixel in an image to its corresponding semantic category, focusing on the geometric structure and differentiation of different objects [SS_survey, SS_survey_2, SS_survey_3]. However, robot indoor semantic segmentation faces several challenges, e.g. complex indoor scenes, occlusions, lighting variations, as well as the diversity of object shapes and sizes. Four RGB-D semantic segmentation algorithms, i.e. ACNet [ACNet], RedNet [RedNet], ESANet [ESANet], and SA-Gate [SA-Gate], are selected and tested on the aforementioned synthetic supermarket scene and real-world canteen scene. 31 and 19 static and dynamic objects with semantic labels are added to the supermarket and canteen scene, respectively. Objects without labels are set to the void value with corresponding pixels (0, 0, 0).
The accuracy of different semantic segmentation methods is measured by calculating the MIoU (Mean Intersection over Union) during testing and the results are shown in Table III. We can obtain three conclusions. First, the accuracy on real-world canteen scene is much lower than synthetic supermarket scene. Second, the MIoU results are comparable to their original published results, though tested on different datasets. Third, to obtain a quantitative comparison on dynamic objects, we also train and test the ESANet [ESANet] algorithm in the supermarket scene with and without dynamic objects, the MIoU of the test is 78.42% and 79.63% respectively. There is no significant difference, which is coincident with our intuition. In addition, as shown in Fig. 7, different methods still have varying segmentation ability on dynamic object, such as the people and shopping carts.
@@ IV-C Robot Relocalization
Robot relocalization refers to the process of accurately determines and adjusts its position based on the environmental perception information and its own localization capabilities, during its movement [posenet, torii2013visual, dfnet]. Existing methods can be coarsely classified into global- and local-feature based methods.
We conducted extensive experiments on THUD using a representative global- and local-feature based methods, i.e. NetVlad [NetVLAD] and FeatLoc [FeatLoc], respectively. We randomly selected two sequences as testing and the results on both synthetic and real-world scene are visualized in Fig. 6.
In addition, to analyze the effect of the number of dynamic objects (primarily pedestrians) in various scene images, we define a metric to represent a scene’s dynamic complexity, computed as the average number of dynamic pedestrians per frame across the entire dataset. As shown in Fig. 8, with the scene’s dynamic complexity increasing, the methods’ accuracy decreases and FeatLoc [FeatLoc] has a higher drop in accuracy compared to NetVlad [NetVLAD], attributed to dynamic objects introducing disturbances and noise in the extraction of scene image features.